# Parameter-Efficient Fine-Tuning with Attributed Patch Semantic Graph for Automated Patch Correctness Assessment

### v1

@@ 5.4 Metrics
  • F1-score: the metric that weighs the accuracy and recall, given by $2*(Precision*Recall)/(Precision+Recall)$.
@@ 5.5 Implementation Details
Our model is implemented using the Pytorch [45] framework. Following the previous APCA work, we use the Adam optimizer [32] to update the model parameters. As the training process progresses, the learning rate is adjusted (ranging from 0 to 0.00005) to adapt to the model learning at different stages. The maximum sequence length is set to be 1024, and words outside the range are ignored. Besides, the dimension of LoRA is set to be 32. Our implementation and evaluation are performed on an Ubuntu 22.04.5 server equipped with two RTX A6000 GPUs.
@@ 6 Experimental Result
We compare our method with the selected baselines in the APCA field using the Wang, Merge, Balance, and Lin datasets, and the results are respectively shown in Table III, Table IV, Table V, Table VI. For the results of the baselines, we get them according to the way described in Section 5.3, and the ‘-’ symbol in the tables indicates that the result has not been reported in the corresponding article. Since both LLM4PatchCorrect and our approach obtain the best result when Llama3 LLM is used, we refer to the results obtained using Llama3 LLM below when mentioning the results of LLM4PatchCorrect and our approach. However, note that the results obtained using the other two LLMs (CodeLlama and StarCoder) show similar trend.
Table III shows the results obtained for the Wang dataset. From the results, we can see that our method outperforms all static methods using the four metrics on the Wang dataset. In particular, compared with the APPT method (the state-of-the-art machine learning based method), our method is 5.1%, 4.6%, 2.5%, and 3.7% higher in accuracy, precision, recall, and F1 score, respectively. Compared with the LLM4PatchCorrect method (the most advanced LLM based method), our method is 3.0%, 2.9%, 1.5%, and 2.3% higher in accuracy, precision, recall, and F1 score, respectively. Among the dynamic methods, Opad relies on a large number of generated test cases to achieve better results in precision and it takes a lot of time to assess patches. Currently, our method is the closest to Opad among static methods, and it significantly outperforms all dynamic methods in comprehensive evaluation metrics such as F1 score. This result suggests that our method better captures important information for patch correctness prediction and improves the performance of LLMs on the APCA task.
TABLE III: Effectiveness comparison on the Wang dataset.
@@
R-Opad 37.7% 100.0% 14.7% 25.6%
Static-traditional S3 69.7% 79.3% 78.9% 79.0%
ssFix 69.2% 78.9% 78.8% 78.8%
CapGen 68.0% 78.3% 77.4% 77.8%
Static-ML ODS 88.9% 90.4% 94.8% 92.5%
Cache 90.8% 92.9% 94.5% 93.7%
APPT 91.3% 93.2% 94.7% 93.9%
Static-LLM
@@
CodeLlama
95.1% 96.4% 96.2% 96.3%
Graph-LoRA-
StarCoder
95.7% 96.6% 95.8% 96.2%
Graph-LoRA-
@@
96.4% 97.8% 97.3% 97.6%
TABLE IV: Effectiveness comparison on the Merge dataset.
@@
Dynamic PATCH-SIM - - - -
E-Opad 90.2% 99.4% 13.8% 24.2%
R-Opad 90.2% 96.5% 16.4% 28.0%
Static-traditional S3 90.2% 90.4% 90.4% 90.4%
ssFix 90.2% 90.1% 90.1% 90.1%
CapGen 90.2% 90.5% 90.5% 90.5%
Static-ML ODS - - - -
@@
CodeLlama
95.7% 95.9% 94.8% 95.3%
Graph-LoRA-
StarCoder
95.9% 96.2% 94.9% 95.5%
Graph-LoRA-
@@
96.8% 97.2% 95.7% 96.4%
Table IV shows the results obtained for the Merge dataset. From table IV, we can see that our method outperforms all baselines in terms of accuracy, recall and F1 score and outperforms all baselines except E-Opad in terms of precision. Compared with the APPT method, our method is 4.6%, 4.7%, 3.6%, and 4.1% higher in accuracy, precision, recall, and F1 score, respectively. Compared with the LLM4PatchCorrect method, our method is 2.3%, 2.1%, 1.5%, and 1.8% higher in accuracy, precision, recall, and F1 score, respectively. This result again proves that our method can achieve excellent performance in the accurate manually labeled dataset.
TABLE V: Effectiveness comparison on the Balance dataset.
@@
Dynamic PATCH-SIM - - - -
E-Opad 58.4% 96.00% 17.71% 29.9%
R-Opad 55.4% 74.58% 16.24% 26.8%
Static-traditional S3 44.2% 44.3% 44.3% 44.3%
@@
CapGen 49.0% 49.1% 49.1% 49.1%
Static-ML ODS - - - -
@@
CodeLlama
81.7% 82.3% 81.0% 81.2%
Graph-LoRA-
StarCoder
81.7% 82.5% 81.4% 81.4%
Graph-LoRA-
@@
85.8% 86.6% 86.4% 86.5%
TABLE VI: Effectiveness comparison on the Lin dataset.
@@
91.2% 91.6% 91.1% 91.4%
Table V shows the results obtained for the Balance dataset. With regard to this dataset, our method still outperforms all baseline methods in terms of accuracy, recall, and F1 score, and is also better than all baseline methods except E-Opad in terms of precision. Notably, our method has a more obvious improvement on the Balance dataset than the Wang dataset. Compared with the APPT method, our method improves accuracy, precision, recall, and F1 score by 14.0%, 13.9%, 12.8%, and 13.4% respectively. Compared with the LLM4PatchCorrect method, our method improves accuracy, precision, recall, and F1 score by 6.6%, 6.5%, 5.6%, and 6.1% respectively. This demonstrates that our method is more suitable for the situation where the number of positive samples and that of negative samples are balanced.
Table VI shows the results obtained for the Lin dataset. Note that the work by Yang et al. [70] does not use this dataset, so we do not have results for some selected baselines. Similarly, we can see from the table that our method outperforms all APCA baseline methods based on machine learning and LLMs. Compared with the APCA method APPT, our method outperforms it by 11.5%, 10.8%, 7.9%, and 9.6% in accuracy, precision, recall, and F1 score, respectively. Compared with the APCA method LLM4PatchCorrec, our method outperforms it by 5.0%, 3.2%, 2.1%, and 2.7% in accuracy, precision, recall, and F1 score respectively. Note that compared with the Wang dataset and the Merge dataset, this dataset is more balanced and our method again has a more obvious improvement.
Answer to RQ1: Overall, our experimental results show that: (1) Our method outperforms all static APCA methods in all metrics and datasets; (2) Compared with the state-of-the-art APCA method LM4PatchCorrect, our method improves accuracy, precision, recall and F1 score by 2.3% to 6.6%, 2.1% to 6.5%, 1.5% to 5.6%, and 1.8% to 6.1% respectively; (3) Our method achieves better performance for the situation where the number of correct patches and that of the overfitting patches are balanced.
@@
To demonstrate the effectiveness of each sub-model and show its contribution to the final results, we perform ablation experiments using the three considered LLMs Llama3, StarCoder, and CodeLlama. In addition, considering whether the number of correct patches and overfitting patches is balanced, we conduct ablation experiments using the Wang dataset and Balance dataset. The former is an imbalanced dataset and the latter is a balanced dataset. We start with the complete model, and then gradually remove specific parts and observe the results after removal. Specifically, to observe the role of the attention mechanism, we first remove the attention fusion layer of Graph-LoRA and replace it with Graph-LoRA-Weak which achieves fusion through vector concatenation; we then remove Graph-LoRA-Weak and directly input the linearized APSG content into the LLM in the form of sequences to observe whether GNNs are more effective in acquiring graph information than linearizing the graph; we next delete the attributes of APSG and only input the graph structure of APSG and code patch into the LLM to observe the role of the patch attributes; we finally remove the whole APSG and only input the code patch into the LLM to observe the role of the graph structure information of APSG. The results obtained are shown in Table VII and Table VIII, and we can have the following observations according to these two tables.
TABLE VII: Ablation Study on the Wang dataset.
@@
Graph-LoRA-Llama3 96.4% 97.8% 97.3% 97.6%
-Graph-LoRA-Attention 95.7% 96.6% 96.1% 96.3%
-Graph-LoRA-Weak 94.2% 94.2% 94.3% 94.2%
-APSG-Attribute 93.1% 93.3% 93.6% 93.4%
-APSG-Graph 91.3% 91.5% 91.9% 91.7%
Graph-LoRA-StarCoder 95.7% 96.6% 95.8% 96.2%
-Graph-LoRA-Attention 94.9% 95.8% 95.1% 95.4%
-Graph-LoRA-Weak 93.4% 93.9% 93.4% 93.6%
-APSG-Attribute 92.5% 92.6% 92.8% 92.7%
-APSG-Graph 90.3% 90.8% 91.1% 90.9%
Graph-LoRA-CodeLlama 95.1% 96.4% 96.2% 96.3%
-Graph-LoRA-Attention 94.2% 95.4% 95.3% 95.3%
-Graph-LoRA-Weak 93.0% 93.6% 93.2% 93.4%
-APSG-Attribute 92.3% 92.4% 92.7% 92.5%
-APSG-Graph 89.8% 90.5% 90.7% 90.6%
TABLE VIII: Ablation Study on the Balance dataset.
@@
Graph-LoRA-Llama3 85.8% 86.6% 86.4% 86.5%
-Graph-LoRA-Attention 84.5% 85.8% 85.2% 85.5%
-Graph-LoRA-Weak 82.2% 83.5% 82.7% 83.1%
-APSG-Attribute 81.4% 82.7% 82.1% 82.4%
-APSG-Graph 78.9% 80.3% 79.6% 79.9%
Graph-LoRA-StarCoder 81.7% 81.5% 81.4% 81.4%
-Graph-LoRA-Attention 80.3% 80.9% 80.1% 80.5%
-Graph-LoRA-Weak 77.5% 78.4% 77.8% 78.1%
-APSG-Attribute 76.8% 77.4% 77.1% 77.2%
-APSG-Graph 74.2% 74.8% 74.3% 74.5%
Graph-LoRA-CodeLlama 81.7% 81.3% 81.0% 81.2%
-Graph-LoRA-Attention 80.1% 80.2% 79.9% 80.0%
-Graph-LoRA-Weak 77.2% 77.5% 77.2% 77.4%
-APSG-Attribute 76.5% 76.8% 76.2% 76.5%
-APSG-Graph 74.2% 74.3% 73.9% 74.1%
@@
First, after removing the attention mechanism within Graph-LoRA, the performance of the model decreases. In the imbalanced Wang dataset, the performance of the model decreases by 0.7% to 0.9%, 0.8% to 1.2%, 0.7% to 1.2%, and 0.8% to 1.3% in terms of the accuracy, precision, recall, and F1 score respectively. In the balanced Balance dataset, the performance of the model in terms of the accuracy, precision, recall, and F1 score decreases by 1.3% to 1.6%, 0.6% to 1.1%, 1.1% to 1.3%, and 0.9% to 1.2% respectively. This shows that compared to directly concatenating graph features and text features, the attention mechanism can better help LLMs acquire graph information.
@@
Second, after removing Graph-LoRA-Weak, we find that the performance of the model decreases significantly. In the imbalanced Wang dataset, the performance of the model decreases by 1.2% to 1.5%, 1.7% to 2.4%, 1.7% to 2.1%, and 1.5% to 2.1% in terms of the accuracy, precision, recall, and F1 score respectively. In the balanced Balance dataset, the performance of the model in terms of the accuracy, precision, recall, and F1 score decreases by 2.3% to 2.9%, 2.3% to 2.7%, 2.3% to 2.7%, and 2.4% to 2.6% respectively. This shows that compared to inputting linearized graph information into LLM, GNNs can obtain graph information more effectively .
@@
Third, after deleting the attribute of APSG and keeping only the graph structure of APSG, the performance of the model also drops. In the imbalanced Wang dataset, the performance of the model drops by 0.8% to 1.1%, 0.9% to 1.3%, 0.5% to 0.7%, and 0.8% to 0.9% in terms of the accuracy, precision, recall, and F1 scores respectively. In the balanced Balance dataset, the performance of the model in terms of the accuracy, precision, recall, and F1 score decreases by 0.7% to 0.8%, 0.7% to 1.0%, 0.6% to 1.0%, and 0.7% to 0.9% respectively. This suggests that the explicit code attributes can help LLM determine the correctness of the patches.
@@
Finally, we delete the whole APSG and keep only the linearized code patch. In the imbalanced Wang dataset, the performance of the model decreases by 1.8% to 2.5%, 1.8% to 1.9%, 1.7% to 2.0% and 1.7% to 1.9% in terms of the accuracy, precision, recall, and F1 score, respectively. In the balanced Balance dataset, the performance of the model with the accuracy, precision, recall, and F1 score decreases by 2.3% to 2.6%, 2.4% to 2.6%, 2.3% to 2.8%, and 2.4% to 2.7% respectively. This shows that the graph structure information of APSG, which captures the patch semantic through data and control flow between program elements, are vital for helping LLM assess the correctness of the patches.
Answer to RQ2: The performance of the model after removing different sub-modules shows that: (1) all major ingredients of the proposed method contribute positively to the final results; (2) the graph part of APSG and the GNN part of Graph-LoRA have more obvious impact on the results.
@@
TABLE IX: Effectiveness of Graph-LoRA in a cross-project setting on the Wang dataset.
@@
Our 91.7% 92.3% 92.5% 92.4%
TABLE X: Effectiveness of Graph-LoRA in a cross-project setting on the Merge dataset.
@@
Our 92.8% 93.4% 92.7% 93.0%
TABLE XI: Effectiveness of Graph-LoRA in a cross-project setting on the Balance dataset.

### v2

@@ 5.4 Metrics
  • F1-score: the metric that weighs the accuracy and recall, given by 2(PrecisionRecall)/(Precision+Recall)2*(Precision*Recall)/(Precision+Recall).
@@ 5.5 Implementation Details
Our model is implemented using the Pytorch [pytorch] framework. Following the previous APCA work, we use the Adam optimizer [kingma2014adam] to update the model parameters. As the training process progresses, the learning rate is adjusted (ranging from 0 to 0.00005) to adapt to the model learning at different stages. The maximum sequence length is set to be 1024, and words outside the range are ignored. Besides, the dimension of low-rank matrices in Graph-LoRA are set to be 256. Our implementation and evaluation are performed on an Ubuntu 22.04.5 server equipped with two RTX A6000 GPUs.
@@ 6 Experimental Result
We compare our method with the selected baselines in the APCA field using the Wang, Merge, Balance, Lin, and Multi-Benchmarks datasets, and the results are shown in Table IV, Table V, Table VI, Table VII, and Table VIII respectively. For the results of the baselines, we get them according to the way described in Section 5.3 and the ‘-’ symbol in the tables indicates that the result has not been reported in previous works and we also do not reproduce the corresponding APCA method for reasons given in Section 5.3. Since both LLM4PatchCorrect and our approach obtain the best result when Llama3 LLM is used, we refer to the results obtained using Llama3 LLM below when mentioning the results of LLM4PatchCorrect and our approach. However, note that the results obtained using the other two LLMs (CodeLlama and StarCoder) show similar trend.
Table IV shows the results obtained for the Wang dataset. From the results, we can see that our method outperforms all static methods using the four metrics on the Wang dataset. In particular, compared with the APPT method (the state-of-the-art ML-based method), our method is 6.4%, 6.9%, 2.2%, and 4.7% higher in accuracy, precision, recall, and F1 score respectively. Compared with the LLM4PatchCorrect method (the most advanced LLM-based method), our method is 3.4%, 3.5%, 2.5%, and 3.0% higher in accuracy, precision, recall, and F1 score respectively. Among the dynamic methods, Opad relies on a large number of generated test cases to achieve better results in precision and it takes a lot of time to assess patches. Currently, our method is the closest to Opad among static methods, and it significantly outperforms all dynamic methods in comprehensive evaluation metrics such as F1 score. This result suggests that our method better captures important information for patch correctness prediction and improves the performance of LLMs on the APCA task.
TABLE IV: Effectiveness comparison on the Wang dataset.
@@
Dynamic R-Opad 37.7% 100.0% 14.7% 25.6%
S3 69.6% 79.1% 79.1% 79.1%
ssFix 69.2% 78.8% 78.8% 78.8%
Static- traditional CapGen 68.1% 78.0% 78.0% 78.0%
ODS 88.9% 90.4% 94.8% 92.5%
BERT_LR 89.4% 90.7% 95.1% 92.5%
BATS 89.8% 90.8% 95.6% 93.1%
PANTHER 91.0% 91.3% 95.3% 93.3%
Cache 90.1% 91.2% 94.5% 92.9%
Static-ML APPT 90.4% 91.5% 96.0% 93.6%
@@
Graph-LoRA-
CodeLlama
95.8% 96.9% 97.6% 97.3%
Graph-LoRA-
StarCoder
96.1% 96.9% 97.8% 97.4%
@@
96.8% 98.4% 98.2% 98.3%
TABLE V: Effectiveness comparison on the Merge dataset.
@@
PATCH-SIM - - - -
E-Opad 22.2% 99.4% 13.8% 24.2%
Dynamic R-Opad 24.1% 96.5% 16.4% 28.0%
S3 82.6% 90.4% 90.4% 90.4%
ssFix 82.1% 90.1% 90.1% 90.1%
Static- traditional CapGen 82.9% 90.5% 90.5% 90.5%
ODS - - - -
BERT_LR 90.9% 91.3% 91.0% 91.1%
BATS 91.4% 91.2% 91.4% 91.3%
PANTHER 92.1% 92.1% 91.7% 91.9%
@@
Graph-LoRA-
CodeLlama
96.5% 96.8% 96.6% 96.7%
Graph-LoRA-
StarCoder
96.7% 96.8% 96.6% 96.7%
@@
97.6% 97.6% 97.9% 97.8%
Table V shows the results obtained for the Merge dataset. From table V, we can see that our method outperforms all baselines in terms of accuracy, recall and F1 score and outperforms all baselines except E-Opad in terms of precision. Compared with the APPT method, our method is 5.4%, 5.1%, 5.8%, and 5.5% higher in accuracy, precision, recall, and F1 score respectively. Compared with the LLM4PatchCorrect method, our method is 3.1%, 2.5%, 3.7%, and 3.2% higher in accuracy, precision, recall, and F1 score respectively. This result again proves that our method can achieve excellent performance in the accurate manually labeled dataset.
TABLE VI: Effectiveness comparison on the Balance dataset.
@@
PATCH-SIM - - - -
E-Opad 58.5% 96.0% 17.7% 29.9%
Dynamic R-Opad 55.4% 74.6% 16.2% 26.7%
S3 44.3% 44.3% 44.3% 44.3%
@@
Static- traditional CapGen 49.1% 49.1% 49.1% 49.1%
ODS - - - -
BERT_LR 63.4% 61.7% 63.4% 62.5%
BATS 64.2% 63.2% 64.7% 63.9%
PANTHER 69.5% 67.6% 69.7% 68.6%
@@
Graph-LoRA-
CodeLlama
82.5% 83.8% 82.6% 83.2%
Graph-LoRA-
StarCoder
82.8% 84.2% 83.1% 83.6%
@@
86.7% 87.8% 87.2% 87.5%
TABLE VII: Effectiveness comparison on the Lin dataset.
@@
Graph-LoRA-
CodeLlama
90.7% 90.7% 90.2% 90.5%
Graph-LoRA-
StarCoder
90.7% 90.4% 89.8% 90.1%
Our
Graph-LoRA-
Llama3
91.8% 92.5% 92.2% 92.3%
Table VI shows the results obtained for the Balance dataset. With regard to this dataset, our method still outperforms all baseline methods in terms of accuracy, recall, and F1 score, and is also better than all baseline methods except E-Opad in terms of precision. Notably, our method has a more obvious improvement on the Balance dataset than the Wang dataset. Compared with the APPT method, our method improves accuracy, precision, recall, and F1 score by 15.7%, 15.1%, 13.6%, and 14.4% respectively. Compared with the LLM4PatchCorrect method, our method improves accuracy, precision, recall, and F1 score by 7.5%, 7.7%, 6.4%, and 7.1% respectively. This demonstrates that our method is more suitable for the situation where the number of positive samples and that of negative samples are balanced.
TABLE VIII: Effectiveness comparison on the Multi-Benchmarks dataset.
Category Method Accuracy Precision Recall F1
Static-ML BERT_LR 80.6% 79.6% 83.3% 81.4%
BATS 82.1% 82.4% 84.2% 83.3%
PANTHER 85.0% 84.2% 85.0% 84.6%
CACHE 82.8% 82.0% 84.1% 83.0%
APPT 85.6% 84.8% 85.5% 85.1%
Static-LLM
LLM4PatchCorrect-
CodeLlama
89.0% 88.2% 89.5% 88.8%
LLM4PatchCorrect-
StarCoder
89.4% 88.8% 90.0% 89.4%
LLM4PatchCorrect-
Llama3
90.2% 89.4% 90.6% 90.0%
@@
94.6% 93.3% 94.8% 94.0%
Table VII shows the results obtained for the Lin dataset. Note that the work by Yang et al. [yang2023large] does not use this dataset, so we do not have results for some selected baselines. Similarly, we can see from the table that our method outperforms all APCA baseline methods based on machine learning and LLMs. Compared with the APPT method, our method outperforms it by 12.6%, 11.7%, 9.0%, and 10.5% in accuracy, precision, recall, and F1 score respectively. Compared with the LLM4PatchCorrect method, our method outperforms it by 6.1%, 4.1%, 3.2%, and 3.6% in accuracy, precision, recall, and F1 score respectively. Note that compared with the Wang dataset and the Merge dataset, this dataset is more balanced and our method again has a more obvious improvement.
Table VIII shows the results obtained for the newly constructed Multi-Benchmarks dataset. It can be seen from the table that our method outperforms all baselines, achieving 94.6%, 93.3%, 94.5%, and 93.0% in terms of accuracy, precision, recall, and F1 score respectively. Compared with the APPT method, our method outperforms it by 9.0%, 8.8%, 9.0%, and 8.8% in accuracy, precision, recall, and F1 score respectively. Compared with the LLM4PatchCorrect method, our method outperforms it by 4.4%, 3.9%, 4.2%, and 4.0% in accuracy, precision, recall, and F1 score respectively. Overall, the result again demonstrates that our method can achieve that best performance on the APCA task.
Answer to RQ1: Overall, our experimental results show that: (1) Our method outperforms all static APCA methods in all metrics and datasets; (2) Compared with the state-of-the-art APCA method LM4PatchCorrect, our method improves the accuracy, precision, recall and F1 score by 3.1% to 7.5%, 2.5% to 7.7%, 2.6% to 6.4%, and 3.0% to 7.1% respectively; (3) Our method achieves better performance for the situation where the number of correct patches and that of the overfitting patches are balanced.
@@
To demonstrate the effectiveness of each sub-element and show its contribution to the final results, we perform two ablation studies using the three considered LLMs Llama3, StarCoder, and CodeLlama. In addition, considering the ratio between the number of correct patches and the number of overfitting patches, the ablation studies are conducted using the Wang dataset (imbalanced) and the Balance dataset (balanced).
The first ablation study focuses on the model structure to demonstrate the contributions of the sub-modules in the model to the final results, and the second ablation study focuses on APSG nodes to demonstrate the contributions of different nodes in APSG to the final results. For the first ablation study, we start with the complete model and then gradually remove specific components and observe the results after removal. Specifically, to observe the role of the attention mechanism, we first remove the attention fusion layer of Graph-LoRA and replace it with Graph-LoRA-Weak which achieves fusion through vector concatenation. To observe whether GNNs are more effective in acquiring graph information than linearizing the graph, we then remove Graph-LoRA-Weak and directly input the linearized APSG content into the LLM in the form of sequences. Next, we delete the attributes of APSG and only input the graph structure of APSG and code patch into the LLM to observe the role of the patch attributes. After that, we remove the whole APSG and only input the code patch into the LLM to observe the role of the graph structure information of APSG. Finally, we do not train the model and only give the LLMs the prompt ”You are a model responsible for assessing patch correctness. Assess whether the patch is correct.” to prove the effectiveness of training. For the second ablation study, we gradually remove the variable nodes, control nodes, and context nodes from the complete model and observe the experimental results. Since APSG is built around patches, we do not remove the patch nodes to preserve the meaning of APSG. The obtained results for the first ablation study are shown in Tables IX and X, and the obtained results for the second ablation study are shown in Tables XI and XII.
TABLE IX: Ablation Study for model structure on the Wang dataset.
@@
Graph-LoRA-Llama3 97.3% 98.6% 98.4% 98.5%
-Graph-LoRA-Attention 96.6% 97.6% 97.2% 97.4%
-Graph-LoRA-Weak 95.1% 95.1% 95.4% 95.3%
-APSG-Attribute 94.1% 94.3% 94.5% 94.4%
-APSG-Graph 92.7% 92.2% 92.0% 92.1%
-Llama3-Train 30.8% 14.3% 35.6% 20.4%
Graph-LoRA-StarCoder 96.5% 97.5% 98.2% 97.8%
-Graph-LoRA-Attention 95.2% 96.4% 97.0% 96.6%
-Graph-LoRA-Weak 93.6% 94.3% 94.5% 94.4%
-APSG-Attribute 92.7% 93.2% 93.6% 93.4%
-APSG-Graph 90.8% 91.0% 91.2% 91.1%
-StarCoder-Train 28.3% 10.7% 2.4% 3.9%
Graph-LoRA-CodeLlama 96.2% 97.2% 97.8% 97.5%
-Graph-LoRA-Attention 95.3% 96.2% 96.8% 96.5%
-Graph-LoRA-Weak 93.6% 94.2% 94.4% 94.3%
-APSG-Attribute 92.7% 93.0% 93.3% 93.2%
-APSG-Graph 90.6% 90.2% 90.5% 90.4%
-CodeLlama-Train 17.6% 1.3% 1.7% 1.5%
TABLE X: Ablation Study for model structure on the Balance dataset.
@@
Graph-LoRA-Llama3 86.7% 87.8% 87.2% 87.5%
-Graph-LoRA-Attention 85.5% 86.8% 86.1% 86.4%
-Graph-LoRA-Weak 83.2% 84.1% 83.6% 83.8%
-APSG-Attribute 82.3% 83.5% 82.8% 83.1%
-APSG-Graph 79.7% 81.0% 80.6% 80.8%
-Llama3-Train 21.6% 4.1% 20.7% 6.8%
Graph-LoRA-StarCoder 82.8% 84.2% 83.1% 83.6%
-Graph-LoRA-Attention 81.4% 82.6% 81.8% 82.1%
-Graph-LoRA-Weak 78.7% 79.6% 78.9% 79.2%
-APSG-Attribute 78.0% 78.8% 78.3% 78.4%
-APSG-Graph 75.3% 75.8% 75.3% 75.5%
-StarCoder-Train 18.3% 2.3% 1.5% 1.8%
Graph-LoRA-CodeLlama 82.5% 83.8% 82.6% 83.2%
-Graph-LoRA-Attention 81.0% 82.2% 81.1% 81.6%
-Graph-LoRA-Weak 78.3% 79.2% 78.5% 78.7%
-APSG-Attribute 77.2% 78.1% 77.5% 77.6%
-APSG-Graph 74.9% 75.7% 75.0% 75.2%
-CodeLlama-Train 14.2% 1.6% 1.3% 1.5%
@@
For the first ablation study, we can have the following observations. First, the model performance decreases after removing the attention mechanism within Graph-LoRA. In the imbalanced Wang dataset, the model performance decreases by 0.7% to 1.1%, 0.9% to 1.1%, 1.0% to 1.2%, and 1.0% to 1.2% in terms of the accuracy, precision, recall, and F1 score respectively. In the balanced Balance dataset, the model performance in terms of the accuracy, precision, recall, and F1 score decreases by 1.2% to 1.5%, 1.0% to 1.6%, 1.1% to 1.5%, and 1.1% to 1.6% respectively. This shows that compared to directly concatenating graph features and text features, the attention mechanism can better help LLMs acquire graph information.
@@
Second, after removing Graph-LoRA-Weak, the model performance decreases significantly. In the imbalanced Wang dataset, the model performance decreases by 1.5% to 1.7%, 1.8% to 2.5%, 1.8% to 2.5%, and 2.1% to 2.2% in terms of the accuracy, precision, recall, and F1 score respectively. In the balanced Balance dataset, the model performance in terms of the accuracy, precision, recall, and F1 score decreases by 2.3% to 2.7%, 2.7% to 3.0%, 2.5% to 2.9%, and 2.6% to 2.9% respectively. This shows that compared to inputting linearized graph information into LLM, GNNs can obtain graph information more effectively.
TABLE XI: Ablation Study for APSG nodes on the Wang dataset.
Model Accuracy Precision Recall F1
Graph-LoRA-Llama3 97.3% 98.6% 98.4% 98.5%
-Variable Node 96.2% 97.3% 97.4% 97.3%
-Control Node 95.9% 97.1% 97.0% 97.1%
-Context Node 93.9% 95.2% 94.6% 94.9%
Graph-LoRA-StarCoder 96.5% 97.5% 98.2% 97.8%
-Variable Node 95.3% 96.4% 97.0% 96.7%
-Control Node 95.0% 95.9% 96.6% 96.2%
-Context Node 92.6% 93.2% 93.5% 93.3%
Graph-LoRA-CodeLlama 96.2% 97.2% 97.8% 97.5%
-Variable Node 94.8% 96.0% 96.7% 96.3%
-Control Node 94.3% 95.6% 96.1% 95.8%
-Context Node 92.2% 92.8% 93.0% 92.9%
TABLE XII: Ablation Study for APSG nodes on the Balance dataset.
Model Accuracy Precision Recall F1
Graph-LoRA-Llama3 86.7% 87.8% 87.2% 87.5%
-Variable Node 85.3% 86.3% 85.9% 86.1%
-Control Node 85.1% 86.0% 85.7% 85.8%
-Context Node 83.0% 83.6% 82.7% 83.1%
Graph-LoRA-StarCoder 82.8% 84.2% 83.1% 83.6%
-Variable Node 81.1% 82.4% 81.3% 81.8%
-Control Node 80.5% 81.9% 80.8% 81.3%
-Context Node 77.2% 78.5% 77.8% 78.1%
Graph-LoRA-CodeLlama 82.5% 83.8% 82.6% 83.2%
-Variable Node 80.8% 82.0% 80.9% 81.4%
-Control Node 80.3% 81.6% 80.3% 80.9%
-Context Node 76.8% 78.1% 77.1% 77.6%
@@
Third, after deleting the attribute of APSG and keeping only the graph structure of APSG, the model performance also drops. In the imbalanced Wang dataset, the model performance drops by 0.9% to 1.0%, 0.8% to 1.4%, 0.9% to 1.1%, and 0.9% to 1.1% in terms of the accuracy, precision, recall, and F1 scores respectively. In the balanced Balance dataset, the model performance in terms of the accuracy, precision, recall, and F1 score decreases by 0.7% to 1.1%, 0.6% to 1.1%, 0.6% to 1.0%, and 0.7% to 1.1% respectively. This suggests that explicit code attributes can help LLM determine the correctness of patches.
@@
Fourth, the performance of the model again obviously decreases if the whole APSG is deleted. In the imbalanced Wang dataset, the performance of the model decreases by 1.4% to 2.1%, 2.1% to 2.8%, 2.4% to 2.8% and 2.7% to 2.8% in terms of the accuracy, precision, recall, and F1 score, respectively. In the balanced Balance dataset, the performance of the model in terms of the accuracy, precision, recall, and F1 score decreases by 2.3% to 2.7%, 2.4% to 2.9%, 2.2% to 3.0%, and 2.3% to 2.9% respectively. This shows that the graph structure information of APSG, which captures the patch semantic through data and control flow between program elements, is vital for helping LLM assess the correctness of the patches.
Finally, if the model is given only the LLM prompt ”You are a model responsible for assessing patch correctness. Assess whether the patch is correct.”, the performance of all three LLMs drops significantly on both datasets, especially for the LLMs StarCoder and CodeLlama. This result proves the effectiveness of our training process. At the same time, the low-quality results of the basic LLMs also suggest that the three LLMs we used likely have no data leakage issue on the studied datasets, and our method is the primary reason for the improvement in model performance.
For the second ablation study, the following observations can be made. First, the model performance obviously decreases after removing the variable nodes. In the imbalanced Wang dataset, the model performance decreases by 1.1% to 1.4%, 1.1% to 1.3%, 1.0% to 1.2% and 1.1% to 1.2% in terms of the accuracy, precision, recall, and F1 score respectively. In the balanced Balance dataset, the model performance with respect to the accuracy, precision, recall, and F1 score decreases by 1.4% to 1.7%, 1.5% to 1.8%, 1.3% to 1.8%, and 1.4% to 1.8% respectively. This shows that variable nodes provide important information in assessing patch correctness. Second, the model performance decreases slightly after removing the control nodes. In the imbalanced Wang dataset, the model performance decreases by 0.3% to 0.5%, 0.2% to 0.5%, 0.4% to 0.6% and 0.2% to 0.5% in terms of the accuracy, precision, recall, and F1 score respectively. In the balanced Balance dataset, the model performance in terms of the accuracy, precision, recall, and F1 score decreases by 0.2% to 0.6%, 0.3% to 0.5%, 0.2% to 0.6%, and 0.3% to 0.5% respectively. We believe that the slight decrease in model performance is due to the relatively small proportion of control nodes. Nevertheless, the control nodes still obviously enrich the patch semantic information. Finally, the model performance decreases significantly after removing the context nodes in APSG. In the imbalanced Wang dataset, the model performance decreases by 2.0% to 2.4%, 1.9% to 2.8%, 2.4% to 3.1% and 2.2% to 2.9% in terms of the accuracy, precision, recall, and F1 score respectively. In the balanced Balance dataset, the model performance with respect to the accuracy, precision, recall, and F1 score decreases by 2.1% to 3.5%, 2.4% to 3.5%, 3.0% to 3.2%, and 2.7% to 3.3% respectively. This result demonstrates that the context nodes store vital semantic information and serve as an important basis for assessing patch correctness.
@@
TABLE XIII: Effectiveness of our method in a cross-project setting on the Wang dataset.
@@
Our 92.4% 93.1% 93.3% 93.2%
TABLE XIV: Effectiveness of our method in a cross-project setting on the Merge dataset.
@@
Our 93.4% 94.0% 93.3% 93.6%
Answer to RQ2: The performance of the model after removing different elements in two ablation studies shows that: (1) all the main components of the proposed method contribute positively to the final results, especially the graph part of APSG and the GNN part of Graph-LoRA; (2) different types of nodes in APSG all contribute to the final result to varying degrees, and the context nodes have the most obvious impact.
TABLE XV: Effectiveness of our method in a cross-project setting on the Balance dataset.